please help solve, it's important
x^y + y^x =17.....1
x + y =5
solve for x and y
question 2
x^2=√(x^2 + y^2)
find x and y
question 3
3^x - 2^x =65
x =4 but how
thanks

Found 2 solutions by Fombitz, Alan3354:
Answer by Fombitz(32388)

Question 1
.
.
.
.
.
.
.
Question 2, you only have one equation.
Need two to solve for an intersection.
.
.
.
.
.
.
.
Question 3,
.
.
.
.
.

Answer by Alan3354(69443)

please help solve, it's important
x^y + y^x =17.....1
x + y =5
solve for x and y
================
2^3 + 3^2 = 17 (by observation)
If you sub 5-y for x then iterative methods can be used.
===============
question 2
x^2= sqrt(x^2 + y^2)
find x and y
x = ±1, y = 0 by observation
------------------
question 3
3^x - 2^x =65
x =4
There's no algebraic solution.
Use iteration, eg, Newton's method.
---------------
but how
thanks